Consider the linear transformation T: R3 + R2 defined as T(X1, X2, 23)=(-23, -3 &1 – 23).Write the standard matrix for HoT, where H is the reflection of R2 about the y-axis. ab sin (a) a дх f a 12 ?
Answer #1

T : R^3\rightarrow R^2

as we are defining composition H must be from H : R^2 \rightarrow R^2

H_oT(x)=H(T(x))   x = (x_1,x_2,x_3) \in R^3

H_oT : R^3 \rightarrow R^2

so the standard matrix must be 2 X 3 matrix

H is a reflection from y \, axis

therefore

H\begin{bmatrix} x\\y \end{bmatrix}=\begin{bmatrix} -x\\ y \end{bmatrix}

H_oT(x) = H(T(x))

  \\*= H(-x_3,-3x_1-x_3) \\*=(x_3,-3x_1-x_3)

so standard matrix for H_oT

\\*H_oT(x)= Ax \\*H_oT(x)=\begin{bmatrix} 0 &0 &1 \\ -3&0 &-1 \end{bmatrix}\begin{bmatrix} x_1\\ x_2\\ x_3 \end{bmatrix}

A = \begin{bmatrix} 0 & 0 & 1\\ -3& 0 &-1 \end{bmatrix} standard matrix .